OpenStudy (fools101):

Which is a factor of P(x) = x4 - 3x3 - x2 - 11x - 4? A) (x + 2) B) (x + 1) C) (x + 4) D) (x - 4) Want to make sure i get this one right is it D?

OpenStudy (fools101):

please help!

OpenStudy (fools101):

is it D?!

jimthompson5910 (jim_thompson5910):

let's check if it's D if it was, then P(4) = 0 So P(4) = (4)^4-3(4)^3-(4)^2-11(4)-4 P(4) = 256-3(64)-16-11(4)-4 P(4) = 256-192-16-44-4 P(4) = 0 So because the remainder is 0, x-4 is a factor

OpenStudy (fools101):

so yes it is D!!! : D

jimthompson5910 (jim_thompson5910):

yep the mouth says it all lol

OpenStudy (fools101):

lol hahah ^ ^... thank u!!

jimthompson5910 (jim_thompson5910):

yw